Quatro Co. issues bonds dated January 1, 2019, with a par value of $850,000. The bonds annual contract rate is 12%, and inteQuatro Co. issues bonds dated January 1, 2019, with a par value of $850,000. The bonds annual contract rate is 12%, and inteQuatro Co. issues bonds dated January 1, 2019, with a par value of $850,000. The bonds annual contract rate is 12%, and inte
Answer #1
1
Premium 43131 =893131-850000
2
Total interest expense over life of bonds
6 payments of $ 51000 306000
Par value at maturity 850000
Total repaid 1156000
Less: Amount borrowed 893131
Total bond interest expense 262869
3
Semiannual Interest period end Cash interest paid Bond interest expense Premium amortization Unamortized Premium Carrying value
01/01/2019 43131 893131
06/30/2019 51000 44657 6343 36788 886788
12/31/2019 51000 44339 6661 30127 880127
06/30/2020 51000 44006 6994 23133 873133
12/31/2020 51000 43657 7343 15790 865790
06/30/2021 51000 43290 7710 8080 858080
12/31/2021 51000 42920 8080 0 850000
Total 306000 262869 43131
Workings:
Cash interest paid 51000 =850000*12%*6/12
Bond interest expense = Carrying value X 10% X 6/12
Bond interest expense:
06/30/2019 44657 =893131*10%*6/12
12/31/2019 44339 =886788*10%*6/12
06/30/2020 44006 =880127*10%*6/12
12/31/2020 43657 =873133*10%*6/12
06/30/2021 43290 =865790*10%*6/12
12/31/2021 42920 =858080*10%*6/12